OpenStudy (anonymous):

good morning everyone. can you please help? thank you =) EXPRESS THE AREA of a RIGHT TRIANGLE as a function of the base "b" if the base the triangle is twice its height A( )=( )

OpenStudy (radar):

Given b=2h or h=b/2 A= 1/2 b h The area of a triagnle \[A=b/2 (b/2) \]\[A=b ^{2}/4\]

OpenStudy (anonymous):

thank you radar...i had an idea but didn't know how to fill in A( )=( ) thank you again for your help.

OpenStudy (radar):

No problem. Good luck.